OpenStudy (anonymous):

explaination must!!!! An empty water tank of length , breadth and height 3.3m, 2m and 1m respectively is to be filled with water by a tap of radius 3 cm and the rate of flow of water is 7 m/sec. From this tank water is continuously utilized at the same rate of flow of water by a water tap of cross sectional radius of 2cm. Calculate the time after which tank will be completely filled, of both the taps are opened simultaneously?

OpenStudy (anonymous):

??????????????????

OpenStudy (anonymous):

*

OpenStudy (anonymous):

wht

OpenStudy (shubhamsrg):

area of cross section of water coming /sec from tap of radius 0.03 = 0.03^2 *pi =0.0009pi rate of flow of water vertically is 7 m/sec so volume of water /sec = 7 * 9* pi * 10^-4 = 63 pi* 10^-4 volume of tank = 3.3 *2 *1 = 6.6 area of cross section of water flowing out of radius 0.02 = 0.02 ^2 * pi = 0.0004 pi rate of flow is same so volume flowing out/sec = 28 pi* 10^-4 thus, net volume filling in /sec = (63 -28) pi*10^-4 = 35 pi * 10^-4 total time = 6.6 *10^4/ (35 pi) = 600.2 approx secs am not too sure to tell the truth..please confirm the ans..

OpenStudy (anonymous):

nai hint is the ans is a natural no

OpenStudy (shubhamsrg):

it comes out to be 10 mins approximately ..

OpenStudy (anonymous):

hmmmmmmmmmm

OpenStudy (anonymous):

absolutely right

OpenStudy (anonymous):

may i ask u a Q'n

OpenStudy (shubhamsrg):

you may! :P

OpenStudy (anonymous):

do u have an fb acount

OpenStudy (anonymous):

account

OpenStudy (shubhamsrg):

maybe we should chat further in pvt messages ? ;)

OpenStudy (anonymous):

ur id

OpenStudy (anonymous):

username

OpenStudy (anonymous):

mine is amruta bharsakle

OpenStudy (shubhamsrg):

hmm..alright,,you dont seem to respond there.. everyone is on fb these days! you may search me via shubham.shubhu@gmail.com

OpenStudy (anonymous):

g talk

OpenStudy (anonymous):

do u have skype account

OpenStudy (anonymous):

bt do u have an fb act

OpenStudy (shubhamsrg):

hmm..yep..skype id is shubham.srg facebook name is shubham raj, but if you search for shubham.shubhu@gmail.com, you can find me directly..also the same id for gtalk..

OpenStudy (anonymous):

yup i got come online on skype

OpenStudy (shubhamsrg):

going to have dinner,,later.. see you in 10-15 mins.. and am deleting all my messages here..add me on fb only if you wish..am deleting the message..

OpenStudy (anonymous):

not there

OpenStudy (anonymous):

on fb

OpenStudy (shubhamsrg):

shubham.shubhu@gmail.com you sure you typed it right ?

OpenStudy (anonymous):

fb gtalk ka kaam ho gaya

OpenStudy (shubhamsrg):

i sent you req on fb..

OpenStudy (anonymous):

hmmm i accepted

OpenStudy (shubhamsrg):

hmmm..